QUESTION 35 Use the equation Qd = 5,000 - 15P + 50A + 3Px - 4I, (2,117) (2.7) (15) (2) (3) where Qd = Quantity Demanded, P = Good Price, A = Advertising Expenditures, Px = Price of a Competitive Good, A = Advertising Expenditures, I = Average Monthly Income, and the Standard Errors of the Regression Coefficients are shown in Parentheses. Calculate the t-statistics for each variable and explain what inferences can be drawn from them. If R2 of this equation is 0.25, what inference can be drawn from it?
Answer #1

We use Thumb Rule of 2 to check if the coefficients are statistically significant at 5% significance level.

t-statistics for independent variable P=Regression coefficient/SE of coefficient=(-15)/2.7=-5.56

t-statistics for independent variable A=Regression coefficient/SE of coefficient=50/15=3.33

t-statistics for independent variable Px=Regression coefficient/SE of coefficient=3/2=1.50

t-statistics for independent variable I=SE/Regression coefficient=-4/3=-1.33

Absolute value of t is higher than 2 for P and A we can say these independent variables are statistically significant at 5% significance level.

Absolute value of t is lower than 2 for Px and I, we can these independent variable are not statistically significant at 5% significance level.

R2 is 0.25. It means that 25% of variations Qd are explained by the variations in independent variables (i.e. P, A, Px and I)